# Bridging the Gulf of Envisioning: Cognitive Design Challenges in LLM Interfaces

### v1

@@ Abstract.
Large language models (LLMs) exhibit dynamic capabilities and appear to comprehend complex and ambiguous natural language prompts. However, calibrating LLM interactions is challenging for interface designers and end-users alike. A central issue is our limited grasp of how human cognitive processes begin with a goal and form intentions for executing actions, a blindspot even in established interaction models such as Norman’s gulfs of execution and evaluation. To address this gap, we theorize how end-users ‘envision’ translating their goals into clear intentions and craft prompts to obtain the desired LLM response. We define a process of Envisioning by highlighting three misalignments: (1) knowing whether LLMs can accomplish the task, (2) how to instruct the LLM to do the task, and (3) how to evaluate the success of the LLM’s output in meeting the goal. Finally, we make recommendations to narrow the envisioning gulf in human-LLM interactions.
@@ 1. Introduction
Why do people find prompt-based interfaces such as ChatGPT (OpenAI, 2023a) appealing? One reason is that Large Language Models (LLM) powering such interfaces allow unrestricted text inputs that can be abstract, complex, and vague, and still produce outputs that appear authentic and credible. For instance, the end-user can prompt GPT to write a wedding toast or a Wikipedia article about Chocolate and receive a lengthy, well-structured text in response. However, formal and anecdotal evidence (e.g., (Zamfirescu-Pereira et al., 2023; Kocoń et al., 2023; Kaddour et al., 2023; Agrawala, 2023)) suggests that producing optimal outputs similar to what the user might inherently produce with LLM interfaces remains a challenge. End-users need to engage in a natural language interaction (NLI) to convey specific intentions such as adhering to a persuasive argument structure, personalizing the narrative for intended readers, avoiding particular biases or perspectives, etc. Given the LLM’s output, they need to validate facts, verify relevance to the original prompt, check for unintended biases, and assess the comprehensiveness and coherence of the generated content to determine if it satisfies their goal. Formulating effective prompt inputs and assessing outputs is easier said than done. In this work, we draw from theories across HCI and cognitive science to characterize the nature of these challenges for existing LLM interfaces and formulate an interaction model with recommendations for the design of human-LLM interfaces.
Early implementations of NLI emerged from the need to draw abstractions from complex information structures (i.e., databases) and provide an intuitive alternative to structured queries (Hendrix et al., 1978; Perrault and Grosz, 1988). Consequently, these interfaces are designed to support highly specific retrieval tasks using precise syntax to minimize ambiguity. While contemporary NLI systems, at the surface, strive for linguistic variability and conversation-style interactions, they remain essentially function-specific commands — like locating the closest coffee shop (Følstad and Brandtzæg, 2017), updating specific attributes of vector graphics (Laput et al., 2013), controlling task workflows (Subramonyam et al., 2018), or constructing data charts (Setlur et al., 2016). NLI design recommendations such as discoverability, variability of language, ambiguity management, and adaptability (Copestake and Jones, 1990; Tory and Setlur, 2019; De Souza, 1993; Srinivasan et al., 2019; Winograd, 1972; Searle, 1980) are aimed at rigid interaction paradigms over preset functions with limited support for spontaneous or new functionalities (Nadin, 1988).
In contrast, LLMs predict related words through training on vast amounts of diverse textual data. This enables them to simulate human-like conversations and provide a more adaptable approach to end-user interaction. More importantly, they exhibit dynamic capabilities in that they are able to fulfill a broad range of intentions and generate ad-hoc functionality in response to prompt inputs (e.g., rewriting a complex text to make it understandable to a five-year-old). This results in an interaction model where (1) the operational scope is not restricted to pre-programmed tasks, (2) variations in end-user inputs are allowed with an “anything goes” approach, and (3) outputs are probabilistic predictions from the user’s query. Consequently, end-users struggle with the LLM’s flexibility because (unlike an ATM’s graphical interface) there is no structure or convention provided (or required) for formulating inputs. End users do not possess and are unable to acquire a system mental model of capabilities and limitations and may not know how to assess, refine, or correct the generated outputs. For designers, the shift in LLMs from a task-centric to a more dynamic interaction paradigm creates challenges in designing human-LLM interactions.
In characterizing types of human-computer interfaces, Hutchins et al. (Hutchins et al., 1985) offer a core paradigm for explaining interface design challenges in software systems. As the distance between the end user’s goal and the system interface increases, the interaction cost increases. This creates an “execution gulf,” which refers to the gap between a user’s intentions and the actions they must specify for the system to achieve them, while an “evaluation gulf” occurs due to the gap between the system’s output and the user’s understanding or interpretation of it. The concept of execution gulf is immediately evident for text-based interactions in LLMs because no detectable affordances suggest how to structure user intentions into system actions. Similarly, the generative (and often inconsistent) nature of outputs and lack of transparency in the LLM’s process introduces a significant gulf of evaluation, making it challenging for users to perceive, interpret, and evaluate the response to their original inputs. In other words, these two gulfs primarily capture the effort required to accomplish the goals.
For HCI systems, a fundamental assumption is that end-users have clear intentions and criteria for evaluating success in meeting their goals. But, this is not always true for LLMs. We posit that the user’s intention formation is connected to the interface, and is supported by system mental models and available actions. In the absence of these key elements, LLM interactions introduce a new kind of gulf– the gulf of envisioning – a gap in their cognitive process connecting their goal (e.g., writing a Wikipedia article about Chocolate) and formulation of intentions (e.g., include a section on ethical Cacao Farming) which are described to the LLM to elicit the desired response. Concretely, the gulf of envisioning execution characterizes end-users’ struggles to engage in cognitive task processes to foresee intentions and then formulate intentions in a format matching the actual process employed by the LLM to respond to inputs. The gulf of envisioning includes at least three misalignments: (1) not knowing whether LLMs can accomplish the task – capability gap, (2) how to instruct the LLM to do the task, i.e., prompt engineering – language gap, and (3) how to evaluate the LLM’s output – intentionality gap.
@@
In this paper, we address the missing link between goals and intentions and formulate a new model of interaction for human-LLM interfaces in which the intentions are the actions. Our key contributions include (1) a characterization of the LLM paradox based on cognitive challenges, (2) an updated model of human-machine interaction, including the process of envisioning execution, and (3) a set of design patterns and guidelines for human-LLM interfaces along with an analysis of interfaces for three types of generative tasks.
2. HCI and the LLM Paradox
@@
Figure 1. Norman’s seven-stage model is a valuable tool when designing interfaces. However, while the second stage – forming the intention – has often been overlooked in most HCI research, it is a crucial part of human-LLM interactions.
However, an intriguing question remaining underexplored is how end-users conceive intentions to engage with the interface. In Norman’s seven-stage model (Norman, 1986), interaction consists of (1) Establishing the Goal, (2) Forming the Intention, (3) Specifying the Action Sequence, (4) Executing the Action on the System’s Interface, (5) Perceiving the System’s State as a Response to the Action, (6) Interpreting the State, and (7) Evaluating the System State with respect to the Goals and Iterating until the goal is achieved (see Figure 1). In much of HCI work, stage two – forming the intention – is assumed as given (Hornbæk and Oulasvirta, 2017). For instance, when cutting and pasting a paragraph of text in a word processor or clicking on the ‘Bold’ font button, how much do we know (or care) about the underlying intentions leading a user to execute those actions? We posit that while this gap from goal to intention has been inadvertently bypassed in traditional design approaches, it emerges as a critical challenge that must be addressed in human-LLM interactions. In this section, we aim to delve into the overlooked aspect of intention formation during interactions and postulate its role in LLM-powered interfaces.
@@ 2.1. Defining Intentions
At the highest level, an intention is a situated pursuit of a goal that is attainable through the execution of a process (Jackson, 1995; Bertelsen and Bødker, 2003). Research in cognitive task analysis suggests that intentions are not just spontaneously generated but arise from a foundation of knowledge, thought processes, and goal configurations (Clark et al., 2008). Intentions include aspects such as declarative knowledge (understanding “what” needs to be done), procedural knowledge (knowing the “how-to” of a task), decision points (key moments in reaching a goal where decisions are necessary), and cognitive skills (the mental abilities required to carry out the task). Further, intentions are tied to goals through complex hierarchical structures, and they emerge as the user works towards achieving those goals.
@@ 2.2. Role of Intentions in Interactions
How does a user formulate intentions and then specify them as actions to be executed by a system? Let’s start with the more familiar aspect of this question, which is action specification, before delving into intentions. According to Norman, users have (or rather acquire) in their head, mental models of the target system, that provide them with the “predictive and explanatory powers” to understand interactions (Norman, 2014). We refer to these representations about systems as system mental models. These models primarily comprise knowledge about how the system operates, including its constituent parts and their interrelations, their inherent processes, and their impact on the system output (Carroll and Olson, 1988). While many researchers have characterized the nature of system mental models (e.g.,  (Young, 2014; Bayman and Mayer, 1984; Mayer, 1981; Moray, 1987)), the common purpose pertaining to HCI is to equip users to determine which action to execute by allowing them to mentally simulate the action (Staggers and Norcio, 1993). For instance, imagine that a writer (the user) is about to type a long and detailed section header. Before committing, they run a mental simulation using the system mental model: they envision typing the header in full length, foreseeing it might take up too much space or look cumbersome in the document. Consequently, they may consider changing the header size or shortening the text. This mental rehearsal helps them anticipate potential readability issues or aesthetic concerns and select their action accordingly.
@@
In the above example, intention formation has not been considered. That is, when mentally simulating the typing of the header, how did the user conceive the header text (i.e., input to the system mental model) in the first place? Of course, this pertains to the user’s goal; let us assume it is to produce a Wikipedia article on Chocolate. A specific intention is the user’s decision or plan to type a long, detailed section header in the document, say “Ethical and environmental implications of cacao bean farming in various tropical regions and their socio-economic impact.” The formulation of this intention is tied to the underlying cognitive task processes of writing that are independent of the system. Generally, they concern how people accomplish mental tasks through component processes, such as retrieval from memory, transforming and combining ideas, and using procedures, etc. Specific to writing, such a cognitive task process is formulated by Hayes and Flowers, and it involves three intertwined cognitive stages, including planning, translating, and reviewing (Hayes, 2013). The specifics of this model are less important at this point, and we will discuss cognitive task processes in detail in Section 3. But what is important is to understand that these cognitive task processes support intention formation. Even for a simpler goal, such as setting a wake-up alarm, the end user must mentally determine the time they need to leave for work, the time it might take to get ready, plan their commute, etc. This is all part of the cognitive task process.
@@
Critically, these two stages – formulating intentions through cognitive task processes and specifying actions through system mental model simulations – interact. In order to generate their intentions for a given goal, the user must be engaged in the cognitive processes needed for that goal. To specify the actions for the system to take based on formulated intentions, the user must simulate the system’s mental model to determine how selected actions may influence the execution of their intention. In reality, users constantly adapt their intentions based on both their dynamic cognitive task processes (what they want to accomplish) and their system mental model (what can be accomplished through system actions). In the Wikipedia example, the mental simulation of typing a long header may result in the reformulation of the intention to a shorter title, say “Ethical impacts of Cacao farming.” Thus, interaction requires (1) intention (formulation and refinement), (2) system mental models that allow a cognitive walkthrough of how the intention might play out at the end of the interaction, and (3) the ability to mentally project the results of their intended actions (i.e., anticipate the eventual outcome of the interaction).

### v2

# Bridging the Gulf of Envisioning: Cognitive Challenges in Prompt Based Interactions with LLMs

@@ Abstract.
Large language models (LLMs) exhibit dynamic capabilities and appear to comprehend complex and ambiguous natural language prompts. However, calibrating LLM interactions is challenging for interface designers and end-users alike. A central issue is our limited grasp of how human cognitive processes begin with a goal and form intentions for executing actions, a blindspot even in established interaction models such as Norman’s gulfs of execution and evaluation. To address this gap, we theorize how end-users ‘envision’ translating their goals into clear intentions and craft prompts to obtain the desired LLM response. We define a process of Envisioning by highlighting three misalignments on not knowing: (1) what the task should be, (2) how to instruct the LLM to do the task, and (3) what to expect for the LLM’s output in meeting the goal. Finally, we make recommendations to narrow the gulf of envisioning in human-LLM interactions.
large language models, prompt-based interactions, cognitive psychology
journalyear: 2024copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642754isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing Interaction design theory, concepts and paradigmsccs: Human-centered computing Natural language interfacesccs: Computing methodologies Natural language generation
Figure 1. A conceptual framework for LLM-based Interactions along three dimensions: (1) Intent Specificity, (2) Functional Flexibility, and (3) Output Determinacy. Point (a) indicates conventional interfaces such as a Calculator with pre-determined functionality and affordances for interaction. Point (b) represents conversational LLMs such as ChatGPT with dynamic functionality. Gulf of Envisioning is the challenge users face in formulating prompts to generate high-quality outputs, i.e., point (c).
Dimensions of interactions
Figure 1. A conceptual framework for LLM-based Interactions along three dimensions: (1) Intent Specificity, (2) Functional Flexibility, and (3) Output Determinacy. Point (a) indicates conventional interfaces such as a Calculator with pre-determined functionality and affordances for interaction. Point (b) represents conversational LLMs such as ChatGPT with dynamic functionality. Gulf of Envisioning is the challenge users face in formulating prompts to generate high-quality outputs, i.e., point (c).
@@ 1. Introduction
Large language models (LLMs) such as ChatGPT (OpenAI, 2023a) have demonstrated remarkable capabilities in generating content that is novel, coherent, and contextually relevant. These models can perform a wide array of tasks, from writing comprehensive essays to creating artwork and even producing functional software interfaces, showcasing a high degree of creativity and adaptability. However, they also require careful guidance to ensure the generated content is appropriate and in alignment with human goals and intentions. For instance, if an end-user wishes to leverage an LLM to craft a toast and prompts the LLM with “Write a toast for Taylor”, the output may be incomplete without providing the desired qualities. The human must be more specific about their intentions (such as, “Write a heartwarming toast for my best friend Taylor’s retirement party, about 5 minutes long, include a humorous twist, and wish them well on the golf course”). Formal and anecdotal evidence (e.g., (Zamfirescu-Pereira et al., 2023b; Kocoń et al., 2023; Kaddour et al., 2023; Agrawala, 2023)) suggests that effectively prompting LLMs to produce outputs similar to human-generated content remains challenging. If intentions are expressed too vaguely or lacking specific detail, the LLM may generate responses that are generic, irrelevant, or off-topic (Holtzman et al., 2021; Zamfirescu-Pereira et al., 2023b; Kim et al., 2023). Iterating with an LLM can correct and progressively guide generation, but playing a “20-questions” or “Hot or Cold” guessing game may be inefficient for longer output and lead to a local minima within the solution space (Simon and Newell, 1971). Further, humans show fixation on initial examples that interfere with exploring alternative solutions (Jansson and Smith, 1991; Leahy et al., 2020). In this work, we draw from theories across HCI and cognitive science to characterize the nature of the cognitive challenges for humans in dialogic interactions with intelligent generative agents.
As shown in Figure 1, the shift towards LLM-powered interfaces can be characterized along the following three dimensions: (1) Functional Flexibility, (2) Intent Specificity, and (3) Output Determinacy. The vision for artificial general intelligence (AGI) includes LLMs with a robust theory of mind (ToM) of humans (and vice-versa), and would allow effective collaboration across numerous tasks. While not yet at this level of general intelligence (Morris et al., 2023; Ullman, 2023), current LLMs do exhibit dynamic capabilities in that they are able to fulfill a broad range of tasks and generate ad-hoc functionality in response to prompt inputs (e.g., “rewrite these appliance installation instructions for a five-year-old”). This flexibility contrasts with conventional direct manipulation interfaces with pre-determined functionalities (e.g., a calculator). Even contemporary natural language interfaces, while aiming for linguistic variability and conversation-style interactions, remain essentially function-specific, like locating the closest coffee shop (Følstad and Brandtzæg, 2017), updating specific attributes of vector graphics (Laput et al., 2013), controlling task workflows (Subramonyam et al., 2018), or constructing data charts (Setlur et al., 2016).
Second, LLMs exhibit open-ended generative characteristics through the vast quantity of linguistic patterns and information learned during training. Their capacity to generate diverse responses to a given input underscores their utility in creative and conversational tasks. While LLMs can produce determinate, correct solutions to closed-ended problems, they offer great value by generating many different solutions to an open-ended problem (Newell et al., 1972). As indicated along the Output Determinacy (O) dimension in Figure 1, the LLM’s capacity to generate varied and unexpected output defines its inherent unpredictability. Humans need to provide oversight and guidance by validating facts, verifying relevance, checking for biases, and evaluating output quality. Third, end-users can converse with LLMs through prompts expressing goals and intentions at any level of specificity. As indicated in the Intent Specificity (I) dimension in Figure 1, structure and convention in the input are not imposed by the LLM, leaving end-users uncertain about how to formulate input to improve LLM generation. The new capabilities of LLMs greatly extend human-machine interactions along these three dimensions into new territory through an interaction model where (1) operational scope is not restricted to pre-programmed tasks, (2) all input intentions are allowed with an “anything goes” approach, and (3) outputs are probabilistic rather than determinate.
In this work, we examine the transformative impact of generative AI systems for human-machine interaction, focusing on how this shift from conventional interfaces alters the design and usability of interactions on these three dimensions. Hutchins et al. (Hutchins et al., 1985) offer a model of interface-design-challenges in software systems, including an “execution gulf” between user intentions and system actions and an “evaluation gulf” between system output and user understanding of its genesis. Their general principle is that as the distance between the human’s intentions and the system’s interface increases, the costs of interaction increase. LLMs transform human-machine interaction to substantially reduce this distance, and therefore the costs, by defining interaction as human formulation of intentions through natural language dialogue leading to desired output (Zamfirescu-Pereira et al., 2023b; Ko et al., 2004). LLMs narrow the gulf of execution by eliminating conventional needs for action specification and execution (Hutchins et al., 1985), leaving only intention to the user. However, the gulf of evaluation may increase by challenges to perceive, interpret, and evaluate output (Hutchins et al., 1985) given the LLMs’ probabilistic process.
What are the consequences of these new LLM features enabling success on complex tasks – flexibility in functional scope, variation in intention specificity, and probabilistic processes and outputs – on the nature and costs of human interaction? We suggest this new LLM interaction process poses new challenges for people, which we call, “the gulf of envisioning.” Concretely, the gulf of envisioning characterizes the distance between the human’s initial intentions and their formulation of a prompt that foresees how LLM capabilities and training data can be leveraged to generate high-quality output. Envisioning includes at least three challenges for humans interacting with LLM systems: (1) how to set my goals and intentions such that the LLM can accomplish the task – the capability gap, (2) how to best instruct an LLM about my goals (i.e., prompt engineering) – the instruction gap, and (3) what to expect for the LLM’s output – the intentionality gap.
@@
In this paper, we formulate a new model of interaction for human-LLM interfaces in which intentions are the actions. Our key contributions include (1) a characterization of how transformative LLM natural language interfaces yield both expansive functionality and new challenges in bridging intentions and outcomes; (2) an updated model of human-machine interaction identifying the process of envisioning execution; and (3) a set of design patterns and guidelines for human-LLM interfaces along with an analysis of interfaces for three types of generative tasks.
2. Intentions and Interactions in Conventional Software Systems
@@
Figure 2. Expanded View of Norman’s seven-stage model of interaction. Norman’s seven-stage model (blue blocks) is a valuable tool when designing interfaces. However, while the second stage – forming the intention – has often been overlooked in most HCI research, it is a crucial part of human-LLM interactions. We add the cognitive task and system mental blocks (in green) to illustrate the mechanism for intention formation and its interaction with action specification.
However, an intriguing question remaining underexplored is how end-users conceive intentions when engaging with the interface. In Norman’s seven-stage model (Norman, 1986), interaction consists of (1) Establishing the Goal, (2) Forming the Intention, (3) Specifying the Action Sequence, (4) Executing the Action on the System’s Interface, (5) Perceiving the System’s State as a Response to the Action, (6) Interpreting the State, and (7) Evaluating the System State with respect to the Goals and Iterating until the goal is achieved (see Figure 2). In much of HCI work, stage two – forming the intention – is assumed as given (Hornbæk and Oulasvirta, 2017). For instance, when cutting and pasting a paragraph of text in a word processor or clicking on the ‘Bold’ font button, how much do we know (or care) about the underlying intentions leading a user to execute those actions? We posit that while this gap from goal to intention has been inadvertently bypassed in traditional design approaches, it emerges as a critical challenge that must be addressed in human-LLM interactions. In this section, we explore this overlooked aspect of intention formation during interactions and postulate its role in LLM-powered interfaces.
@@ 2.1. Defining Intentions
At the highest level, an intention is a situated pursuit of a goal that is attainable through the execution of a process (Jackson, 1995; Bertelsen and Bødker, 2003) of a certain sequence of actions conceived as leading towards a goal. An intention is an intermediate cognitive state that translates the abstract desire (goal) into concrete actions. Research in cognitive task analysis suggests that intentions are not just spontaneously generated but arise from a foundation of knowledge, thought processes, and goal configurations (Clark et al., 2008). Intentions include aspects such as declarative knowledge (understanding “what” needs to be done), procedural knowledge (knowing the “how-to” of a task), decision points (key moments in reaching a goal where decisions are necessary), and cognitive skills (the mental abilities required to carry out the task). Further, intentions are tied to goals through complex hierarchical structures, and they emerge as the user works towards achieving those goals. This key cognitive science insight was advanced in seminal works by Karl Lashley  (Lashley et al., 1951) and Miller et al. (Miller et al., 1960) in their focus on the hierarchical structure of nested subroutines in human action, opposing behaviorist conceptions of sequential behavior as a chain of stimulus-response associations, and it is further elaborated in current work in cognitive and computational neuroscience  (Botvinick, 2008).
@@ 2.2. Role of Intentions in Interactions
How does a user formulate intentions and then specify them as actions to be executed by a system? Let’s start with the more familiar aspect of this question, which is action specification, before delving into intentions. According to Norman, users have (or rather acquire) in their mind, mentally represented models of the target system that provide them with the “predictive and explanatory powers” to understand interactions (Norman, 2014). We refer to these representations about systems that help with action specification as system mental models. These models primarily comprise knowledge about how the system operates, including its constituent parts and their interrelations, their inherent processes, and their impact on the system output (Carroll and Olson, 1988). While many researchers have characterized the nature of system mental models (e.g.,  (Young, 2014; Bayman and Mayer, 1984; Mayer, 1981; Moray, 1987)), the common purpose pertaining to HCI is to equip users to determine which action to execute by allowing them to mentally simulate the action (Staggers and Norcio, 1993). For instance, imagine that a writer (the user) is about to type a long and detailed section header. Before committing, they run a mental simulation using the system mental model: they envision typing the header in full length, foreseeing it might take up too much space or look cumbersome in the document. Consequently, they may consider changing the header size or shortening the text. This mental rehearsal helps them anticipate potential readability issues or aesthetic concerns and select their action sequence accordingly.
@@
In the above example, intention formation has not been considered. That is, when mentally simulating the typing of the header, how did the user conceive the header text (i.e., input to the system mental model) in the first place? Of course, this pertains to the user’s goal; let us assume it is to produce a Wikipedia article on Chocolate. A specific intention is the user’s decision or plan to type a long, detailed section header in the document, say “Ethical and environmental implications of cacao bean farming in various tropical regions and their socio-economic impact.” The formulation of this intention is tied to the underlying cognitive task processes of writing that are independent of the technological system. Generally, they concern how people accomplish mental tasks through component processes, such as retrieval from memory, transforming and combining ideas, and using procedures, etc. Specific to writing, such a cognitive task process is formulated by Hayes and Flower, and it involves three intertwined cognitive stages, including planning, translating, and reviewing (Hayes, 2013). The specifics of this model are less important at this point, and we will discuss cognitive task processes in detail in Section 3. But what it is important to understand is that these cognitive task processes support intention formation. Even for a simpler goal, such as setting a wake-up alarm, the end user must mentally determine the time they need to leave for work, the time it might take to get ready, plan their commute, etc. This is all part of the cognitive task processes.
@@
Critically, these two stages – formulating intentions through cognitive task processes and specifying actions through system mental model simulations – interact. In order to generate their intentions for a given goal, the user must be engaged in the cognitive processes needed for that goal. To specify the actions for the system to take based on formulated intentions, the user must simulate the system’s mental model to determine how selected actions may influence the execution of their intention. In reality, users constantly adapt their intentions based on both their dynamic cognitive task processes (what they want to accomplish) and their system mental model (what they believe can be accomplished through system actions). In the Wikipedia example, the mental simulation of typing a long header may result in the reformulation of the intention to a shorter title, say “Ethical impacts of Cacao farming.” Thus, interaction requires (1) intention (formulation and refinement), (2) system mental models that allow a cognitive walkthrough of how the intention might play out at the end of the interaction, and (3) the ability to mentally project the results of their intended actions (i.e., anticipate the eventual outcome of the interaction).